# A Computational Theory and Semi-Supervised Algorithm for Clustering

### v1

@@ 3.2 Two-dimensional Synthetic Data
The next example is a two-dimensional data set composed of eight gaussian distributed clusters at different locations and with standard deviations $[0.6,2,0.2,0.7,3,0.4,0.6,0.6]$ (Figure 6). Added to this are many isolated anomalies, and a relatively small anomalous cluster which could also be considered a newly emerging cluster. The total number of points is $10300$, and a small random sample of a $100$ points from the eight clusters are taken as seeds which is illustrated by Figure 7. The number of seeds selected is such that we have at least $10$ per cluster in order to reliably build out the clusters. These would be expected to be provided by a domain expert and help guide the clustering algorithm to desired solutions. Note that the labels can even be a little noisy since mislabelled examples can be ejected out of the group as anomalies during algorithm execution.
@@
Figure 6: Eight gaussian clusters are shown with differing standard deviations ($0.6,2,0.2,0.7,3,0.4,0.6,0.6$) and locations. Added to the data is a scattering of anomalies and also a small anomalous cluster at coordinates $(11,20)$.
@@
Figure 7: A random sample of points (1% of the data) is shown from the gaussian distributed clusters. These form the seeds of the clustering algorithm.
@@
The clustering result is shown in Figure 8 where the eight clusters are largely identified, some points around the clusters are deemed anomalous, and many isolated points and the small cluster labelled as anomalies. Note the difficulty of the task in that some clusters are close or overlapping, while others are located inside another. Hence, there are arguably misclassifications along bordering points. Figure 9 also shows a close up view of one cluster to show the proposed methods ability to highlight fringe and far lying points as outliers with respect to the cluster’s mass of points. Contrasting the results with that of $k$-means ($k=8$), we see in Figure 10 that it unintuitvely partitions the points such that it does not reflect the desired groupings. Clusters originally labelled 3 and 6 are also grouped together and the small anomalous group of points is deemed to be a cluster. $k$-means also partitions all the data points resulting in anomalies being assigned to some cluster. The results obtained with DBSCAN ($mps=15,eps=1.17$) exhibit interesting characteristics. The algorithm successfully identifies and labels many isolated points as outliers, accurately capturing their anomalous nature. It also considers the small group of points to indeed be a cluster. However, it merges the majority of points into one large cluster, leading to a loss of granularity in the clustering results. It also does not report many of the points around the gaussian clusters as anomalies (as in Figure 9). This behavior aligns with the algorithm’s primary objective being to cluster data rather than detect anomalies.
@@
Figure 9: A close up view of cluster $7$ shows how the proposed clustering algorithm not only detects clusters, but also highlights anomalous points as those being unexpectedly distant from the central mass of points.
@@
Figure 11: Results of the DBSCAN algorithm. Noteable observations are that most of the clusters are reported to be a single large cluster and that many outlying points are highlighted as anomalies.
@@ 3.3 Multi-dimensional Realworld Examples
This section presents the clustering algorithm’s results on two widely-used benchmark datasets: MNIST handwritten digits and the 20 newsgroup corpus. These datasets were selected to showcase different application areas and to be easily interpretable for non-experts. Given the high-dimensionality of the raw data in these domains, dimensionality reduction techniques were employed to enhance the clustering outcomes. Specifically, UMAP [7], a method known for its speed and effectiveness with largely default settings was utilised. The distance metric and the desired number of dimensions to map to were the only explicitly specified parameters: the cosine metric was used to address the datasets’ high dimensionality and the number of dimensions set to values that depend on the size of the data.
@@ 3.3.1 MNIST Handwritten Digits
Figure 12 depicts the application of UMAP to 2 dimensions for visualising the smaller MNIST digits dataset, consisting of 1,797 examples available in sklearn [9]. The plot reveals nine prominent clusters along with several smaller groups scattered throughout. The ground truth labels indicate the presence of ten distinct groups, each corresponding to a specific digit. For the clustering algorithm, a mapping to a 10-dimensional space is utilised, and approximately 6% of the data (107 examples) is randomly sampled to provide around 10 labeled items per group, serving as seeds for the algorithm. The clustering results are shown in Figure 13, exhibiting an accuracy of 0.84 according to the ground truth. However, a closer examination of the results unveils an interesting finding: the algorithm not only successfully captures much of the primary clusters but also assigns points that may be outliers with respect to the clusters to the anomalous group. These anomalous parts of clusters or individual small clusters often represent either uniquely written digits, or slight variations within the same class, such as the digit 1 with or without a horizontal baseline [^1] The interactive plots where points in the graph can be visualised are available in the github project at https://github.com/M-Nassir/clustering.
Figure 12: This illustrates the 2-dimensional UMAP of the smaller MNIST data set obtained from sklearn, using the cosine metric. Each data point is color-coded based on its corresponding class label obtained from the provided labels.
Figure 13: The results of applying the clustering algorithm to the smaller MNIST digits data set obtained from sklearn after performing a UMAP transformation. The actual clustering process involved mapping the data to 10 dimensions. However, for the purpose of visualisation, the results are presented in a 2-dimensional mapping. The accuracy of the clustering according to the ground truth labels is 0.84.
For completeness, Figure 14 shows the algorithm results on the full MNIST digits data set comprising 70,000 examples with UMAP reduction to 20 dimensions due to the much larger data set size. Here an accuracy of 0.92 is achieved according to the ground truth labels by leveraging merely 0.5% (350) of labelled examples as seeds. Notably, despite the substantial expansion in the number of examples, only a modest number of additional samples were required for satisfactory results. Anomalies are also found within different segments of the primary clusters, warranting further investigation to discern whether they signify false positives or alternative representations of a handwritten digit.
Figure 14: The results of applying the clustering algorithm to the full MNIST digits data set after performing a UMAP transformation. The actual clustering process involved mapping the data to 20 dimensions since the data set is much larger. However, for the purpose of visualisation, the results are presented in a 2-dimensional mapping. The accuracy of the clustering according to the ground truth labels is 0.92.
@@
The 20 Newsgroups data set in the scikit-learn library can be used as an example of clustering data in the text document space. It contains over 18,000 news articles that are organised into 20 categories. In this specific example, a subset of 6 categories is used. These were selected for their distinctness, resulting in a total of 5,881 examples.
To prepare the data, a standard pre-processing procedure is followed, including lemmatisation, creating a tf-idf matrix, and removing stopwords with a minimum document frequency of five. Due to the high dimensionality of the data, UMAP is employed for dimension reduction, as it effectively preserves the relationships between examples. The resulting reduction to 2 dimensions is illustrated in Figure 15, where each example is coloured according to its ground truth category label. Notably, there is significant overlap between clusters, particularly in the central regions, and a small cluster consisting of examples from all categories can be observed on the far top-left of the plot at approximate coordinates (-8,9).
Figure 15: A 2-dimensional UMAP plot of the 20 Newsgroup data set is shown. Each point is coloured according to its ground truth category.
For clustering purposes, the data is mapped to 10 dimensions and a 2% random sample of labelled examples is included to guide the algorithm. This approach incorporates user input to achieve desired outcomes. The clustering results are shown in Figure 16 (displayed over the 2-dimensional mapping). The algorithm successfully recovers most of the six main clusters, achieving an overall accuracy of 0.74 according to the ground truth labels. Moreover, the algorithm flags many examples located between clusters as anomalies, as well as the small dense cluster on the far top-left. This again highlights that the algorithm performs a clustering as opposed to a partitioning of the data points. Table 1 provides the top 10 keywords for each of the six clusters, along with the anomalous group. The keywords align well with the category labels, while the anomalous group consists of seemingly random words that do not align well to any specific category.
Figure 16: The results of applying the proposed semi-supervised clustering algorithm on 6 of the 20 Newsgroup corpus data set categories after a UMAP reduction to 10 dimensions. The accuracy of the clustering according to the ground truth labels is 0.74.
@@
Anomalies
helmet, thanks, edu, just, like, list, mail, com, dog, wa
comp.windows.x
use, display, application, program, motif, widget, thanks, file, server, window
rec.motorcycles
rider, road, just, like, riding, dod, motorcycle, ride, wa, bike
rec.sport.hockey
espn, season, playoff, play, year, player, hockey, wa, team, game
sci.crypt
escrow, use, nsa, algorithm, phone, government, encryption, clipper, chip, key
soc.religion.christian
christians, bible, people, sin, christ, christian, church, jesus, wa, god
talk.politics.guns
just, weapon, government, law, did, right, fbi, people, wa, gun

### v2

@@ 3.2 Two-dimensional Synthetic Data
The next example is a two-dimensional dataset, named 2d_gauss, composed of eight random Gaussian distributed clusters at different locations and with standard deviations $[0.6,2,0.2,0.7,3,0.4,0.6,0.6]$ (Figure 6). Added to this are many isolated anomalies, and a relatively small anomalous cluster which could also be considered a newly emerging cluster. The total number of points is $10300$, and a small random sample of approximately $100$ points from the eight clusters are taken as seeds which is illustrated by Figure 7. The number of seeds selected is such that we have at approximately $10$-$30$ per cluster in order to reliably build out the clusters. These would be expected to be provided by a domain expert and help guide the clustering algorithm to desired solutions. Note that the labels can even be noisy since mislabelled examples will be ejected out of the group as anomalies during algorithm execution.
@@
Figure 6: Eight Gaussian clusters are shown with differing standard deviations and locations. Added to the data is a scattering of anomalies and also a small anomalous cluster at coordinates $(11,20)$.
@@
Figure 7: A random sample of points (1% of the data) is shown from the Gaussian distributed clusters. These form the seeds of the proposed clustering algorithm.
@@
The clustering result is shown in Figure 8 where the eight clusters are largely identified, some points around the clusters are deemed anomalous, and many isolated points and the small cluster labelled as anomalies. Note the difficulty of the task in that some clusters are close or overlapping, while others are located inside another. Hence, there are arguably misclassifications along bordering points. Figure 9 also shows a close up view of one cluster to show the proposed methods ability to highlight fringe and far lying points as outliers with respect to the cluster’s mass of points. Contrasting the results with that of $k$-means ($k=8$), we see in Figure 10 that it unintuitvely partitions the points and does not reflect the desired groupings. Clusters originally labelled 3 and 6 are also grouped together and the small anomalous group of points is deemed to be a cluster. $k$-means also partitions all the data points resulting in anomalies being assigned to some cluster. The results obtained with DBSCAN exhibit interesting characteristics. The algorithm successfully identifies and labels many isolated points as outliers, accurately capturing their anomalous nature. It also considers the small group of points to be an anomalous cluster. However, it merges the majority of points into one large cluster, leading to a loss of granularity in the clustering results. It also does not report many of the points around the Gaussian clusters as anomalies (as in Figure 9). This behavior aligns with the algorithm’s primary objective being to cluster data rather than detect anomalies. Constrained $k$-means and COP-KMeans also produce results similar to $k$-means, even though constraints are applied using the seed data; the result of applying COP-KMeans is shown in Figure 12. (We encourage readers to view results of other widely used clustering algorithms in the associated GitHub repository referenced on the first page).
@@
Figure 9: A close up view of cluster $7$ shows how the proposed clustering algorithm not only detects clusters, but also highlights anomalous and fringe points.
@@
Figure 11: Results of the DBSCAN algorithm with default parameters. Most of the clusters are reported to be a single large cluster while many outlying points are highlighted as anomalies.
Figure 12: Results of applying the COP-KMeans algorithm, where even though the labelled data is used to form constraints, it still results in many undesired partitions of the data that are similar to $k$-means.
@@ 3.3 Multi-dimensional Realworld Examples
This section presents the proposed clustering algorithm results on a selection of widely used benchmark datasets, alongside the synthetic datasets introduced in the previous section. These datasets were selected for their popularity, varied structural characteristics and differing levels of complexity [12]. All datasets were obtained from the UCI library of machine learning datasets [14] and the scikit-learn library [23]. Evaluation was conducted quantitatively using a range of standard external validation metrics—Purity, V–Measure, Normalised Mutual Information (NMI), Adjusted Rand Index (ARI) and the Fowlkes–Mallows Index (FMI)—–to provide a comprehensive assessment of clustering quality. To focus solely on clustering performance, anomalous points were excluded from these analyses.
Table 1 summarises the main characteristics of the datasets, showing the number of examples (n), features (Dim), and classes. In the yeast and shuttle datasets, minority classes were treated as anomalies due to their incoherent structure and resemblance to noise. These were reduced from $10$ and $7$ classes to $4$ and $3$, respectively. For datasets with more than $30$ features, dimensionality was reduced to $10$ using UMAP [21], following a rule-of-thumb heuristic. This included ionosphere ($34$) and MNIST digits ($64$). An exception was made for cover_type, which was kept in its original high-dimensional form due to computational limitations, but as a consequence serves as a representative large-scale example. In the case of the Newsgroups dataset, which consists of text data, the feature count reflects the dimensionality after lemmatisation, TF-IDF vectorisation, and UMAP reduction. The table also states the percentage of labelled data used for semi-supervised methods; smaller datasets generally required a higher proportion of labels to guide clustering effectively as indicated by the average number of labels per cluster column. All datasets are available in the accompanying GitHub repository. Experiments were run on a MacBook M1 Pro with 16 GB of RAM.
Table 1: Summary of benchmark datasets used in the evaluation.
Dataset n Dim #Classes % Labelled #Lbl/Clust.
1d_gauss 17,528 1 3 0.2 12
2d_gauss 10,300 2 8 1 13
6Newsgroups_UMAP10 10,496 10 6 1 17
MNIST_UMAP10 1,797 10 10 5 9
banknote 1,372 4 2 2 13
breast_cancer 569 30 2 7 20
cover_type 581,012 54 7 0.02 15
glass 214 9 6 30 10
ionosphere_UMAP10 351 10 2 10 17
iris 150 4 3 20 10
land_mines 338 3 5 30 20
pendigits 10,992 16 10 2.5 27
seeds 210 7 3 20 14
shuttle 43,500 9 3 0.2 29
wine 178 13 3 30 17
yeast 1,484 8 4 5 18
Figures 13 and 14 plot the purity and NMI performance of the proposed method against the number of seeds per cluster at initialisation across the datasets averaged over ten runs. They indicate that rather than the number of seeds required to be a percentage of the dataset, generally $10-30$ seeds per cluster are sufficient for effective clustering.
Figure 13: Purity scores of the proposed method as a function of the number of points per cluster.
Figure 14: Normalized Mutual Information (NMI) scores of the proposed method as a function of the number of points per cluster.
Tables 27 present the performance of the proposed semi-supervised clustering method across a diverse suite of benchmark datasets, in comparison with several established unsupervised and semi-supervised approaches. The method consistently delivers strong results across synthetic, real-world, and high-dimensional settings. It ranks among the top performers on structured synthetic datasets such as 1d_gauss and 2d_gauss, validating its correctness in controlled conditions. On more complex benchmarks—such as MNIST_UMAP10 and 6Newsgroups_UMAP10—it achieves the highest scores across all evaluation metrics, outperforming traditional, density-based, and deep clustering baselines, including DEC. While its performance is somewhat lower on certain datasets such as iris, seeds, and wine, where methods like GMM tend to excel, the proposed method remains highly competitive overall. As expected, no single method performs best across all datasets and metrics, but the proposed approach demonstrates a favourable balance of robustness and versatility. (Dashed lines indicate instances where computational constraints prevented clustering.)
Importantly, the gains in clustering quality are achieved with competitive computational cost and it was able to cluster all datasets. It runs efficiently on most datasets, requiring no parameter tuning and only a small fraction of labelled data—underscoring its robustness and ease of deployment. Although runtimes are relatively higher on datasets such as cover_type, MNIST_UMAP10, and pendigits compared to lightweight methods like $k$-means, profiling reveals that the overhead arises primarily from redundant cluster expansion steps and preprocessing in the underlying anomaly detection model. These findings point to clear opportunities for optimisation—both at the algorithmic level and through improvements in code efficiency or implementation language.
In summary, the proposed method strikes a strong balance between accuracy, efficiency, and scalability. Its ability to work with minimal supervision, reject noisy seeds during cluster expansion, and leave outliers or unknown groups unassigned for subsequent analysis further underscores its robustness and versatility—making it a practical and competitive choice for semi-supervised clustering in both research and applied contexts.
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 1.00 1.00 1.00 0.93 1.00 1.00 1.00 1.00 1.00
2d_gauss 0.81 0.88 0.84 0.81 0.77 0.78 0.52 0.84 0.89
6NewsgroupsUMAP10 0.74 0.74 0.78 0.74 0.74 0.72 0.34 0.74 0.89
MNIST_UMAP10 0.88 0.88 0.88 0.87 0.87 0.88 0.97 0.88 0.99
banknote 0.61 0.56 0.61 0.61 0.62 0.56 0.70 0.56 0.62
breast_cancer 0.85 0.95 0.85 0.87 0.86 0.78 0.63 0.85 0.90
cover_type 0.49 0.57 0.49 0.49 0.50
glass 0.59 0.56 0.57 0.62 0.62 0.54 0.49 0.43 0.56
ionosphere_UMAP10 0.70 0.69 0.70 0.70 0.72 0.70 0.64 0.68 0.85
iris 0.89 0.97 0.89 0.89 0.79 0.89 0.69 0.93 0.89
land_mines 0.23 0.38 0.30 0.30 0.45 0.28 0.21 0.26 0.41
pendigits 0.75 0.71 0.77 0.77 0.71 0.72 0.10 0.75 0.69
seeds 0.90 0.93 0.90 0.90 0.91 0.89 0.53 0.86 0.90
shuttle 0.79 0.79 0.83 0.79 0.79 0.79 0.84
wine 0.70 0.85 0.70 0.70 0.78 0.70 0.40 0.70 0.70
yeast 0.49 0.44 0.50 0.49 0.49 0.47 0.36 0.47 0.44
Table 2: Purity metric results
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 1.00 1.00 1.00 0.80 1.00 1.00 1.00 1.00 1.00
2d_gauss 0.87 0.92 0.83 0.83 0.82 0.85 0.72 0.85 0.84
6NewsgroupsUMAP10 0.61 0.61 0.62 0.60 0.60 0.61 0.29 0.62 0.76
MNIST_UMAP10 0.90 0.90 0.90 0.87 0.89 0.90 0.93 0.90 0.98
banknote 0.03 0.01 0.03 0.03 0.03 0.00 0.17 0.01 0.01
breast_cancer 0.46 0.71 0.46 0.50 0.48 0.32 0.00 0.46 0.49
cover_type 0.07 0.17 0.07 0.04 0.11
glass 0.43 0.40 0.40 0.36 0.45 0.39 0.31 0.14 0.36
ionosphere_UMAP10 0.12 0.11 0.12 0.12 0.16 0.12 0.10 0.11 0.42
iris 0.74 0.90 0.74 0.76 0.63 0.77 0.60 0.81 0.77
land_mines 0.01 0.20 0.10 0.12 0.18 0.08 0.00 0.01 0.20
pendigits 0.69 0.68 0.69 0.69 0.68 0.73 0.00 0.74 0.64
seeds 0.69 0.76 0.69 0.69 0.73 0.73 0.29 0.65 0.69
shuttle 0.00 0.00 0.26 0.16 0.00 0.08 0.21
wine 0.43 0.58 0.43 0.43 0.50 0.42 0.00 0.43 0.41
yeast 0.12 0.11 0.16 0.12 0.13 0.13 0.00 0.09 0.08
Table 3: V-Measure metric results
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 1.00 1.00 1.00 0.80 1.00 1.00 1.00 1.00 1.00
2d_gauss 0.87 0.92 0.83 0.83 0.82 0.85 0.72 0.85 0.84
6NewsgroupsUMAP10 0.61 0.61 0.62 0.60 0.60 0.61 0.29 0.62 0.76
MNIST_UMAP10 0.90 0.90 0.90 0.87 0.89 0.90 0.93 0.90 0.98
banknote 0.03 0.01 0.03 0.03 0.03 0.00 0.17 0.01 0.01
breast_cancer 0.46 0.71 0.46 0.50 0.48 0.32 0.00 0.46 0.49
cover_type 0.07 0.17 0.07 0.04 0.11
glass 0.43 0.40 0.40 0.36 0.45 0.39 0.31 0.14 0.36
ionosphere_UMAP10 0.12 0.11 0.12 0.12 0.16 0.12 0.10 0.11 0.42
iris 0.74 0.90 0.74 0.76 0.63 0.77 0.60 0.81 0.77
land_mines 0.01 0.20 0.10 0.12 0.18 0.08 0.00 0.01 0.20
pendigits 0.69 0.68 0.69 0.69 0.68 0.73 0.00 0.74 0.64
seeds 0.69 0.76 0.69 0.69 0.73 0.73 0.29 0.65 0.69
shuttle 0.00 0.00 0.26 0.16 0.00 0.08 0.21
wine 0.43 0.58 0.43 0.43 0.50 0.42 0.00 0.43 0.41
yeast 0.12 0.11 0.16 0.12 0.13 0.13 0.00 0.09 0.08
Table 4: NMI metric results
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 1.00 1.00 1.00 0.77 1.00 1.00 1.00 1.00 1.00
2d_gauss 0.76 0.85 0.73 0.73 0.70 0.72 0.43 0.80 0.78
6NewsgroupsUMAP10 0.61 0.60 0.63 0.60 0.58 0.55 0.12 0.60 0.77
MNIST_UMAP10 0.81 0.81 0.82 0.79 0.81 0.81 0.91 0.82 0.98
banknote 0.05 0.00 0.05 0.05 0.05 0.00 0.02 0.01 0.02
breast_cancer 0.49 0.81 0.49 0.54 0.51 0.29 0.00 0.49 0.62
cover_type -0.00 0.09 -0.00 0.01 0.05
glass 0.27 0.25 0.27 0.22 0.27 0.26 0.16 0.06 0.18
ionosphere_UMAP10 0.15 0.14 0.15 0.15 0.20 0.15 -0.07 0.13 0.49
iris 0.72 0.90 0.72 0.73 0.57 0.73 0.52 0.82 0.74
land_mines -0.01 0.07 0.03 0.04 0.11 0.01 0.00 0.00 0.05
pendigits 0.58 0.54 0.60 0.60 0.56 0.55 0.00 0.60 0.52
seeds 0.72 0.80 0.72 0.72 0.76 0.71 0.05 0.64 0.71
shuttle 0.00 0.00 0.46 0.10 0.00 0.04 0.23
wine 0.37 0.61 0.37 0.37 0.48 0.37 0.00 0.37 0.36
yeast 0.09 0.08 0.11 0.10 0.10 0.06 -0.00 0.07 0.06
Table 5: ARI metric results
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 1.00 1.00 1.00 0.87 1.00 1.00 1.00 1.00 1.00
2d_gauss 0.80 0.87 0.77 0.77 0.75 0.77 0.60 0.82 0.81
6NewsgroupsUMAP10 0.68 0.68 0.70 0.67 0.66 0.64 0.44 0.68 0.81
MNIST_UMAP10 0.84 0.83 0.84 0.81 0.83 0.83 0.92 0.85 0.98
banknote 0.55 0.51 0.55 0.55 0.55 0.50 0.51 0.52 0.60
breast_cancer 0.79 0.91 0.79 0.81 0.80 0.74 0.73 0.79 0.84
cover_type 0.26 0.36 0.26 0.24 0.34
glass 0.51 0.48 0.49 0.40 0.45 0.51 0.43 0.32 0.40
ionosphere_UMAP10 0.59 0.58 0.59 0.59 0.61 0.59 0.54 0.58 0.77
iris 0.81 0.94 0.81 0.82 0.71 0.82 0.71 0.88 0.82
land_mines 0.19 0.31 0.23 0.24 0.29 0.23 0.44 0.20 0.35
pendigits 0.62 0.59 0.64 0.64 0.61 0.61 0.32 0.65 0.57
seeds 0.81 0.87 0.81 0.81 0.84 0.81 0.47 0.76 0.81
shuttle 0.80 0.81 0.84 0.61 0.80 0.80 0.73
wine 0.58 0.74 0.58 0.58 0.66 0.58 0.58 0.58 0.59
yeast 0.39 0.45 0.46 0.39 0.39 0.47 0.53 0.34 0.37
Table 6: FMI metric results
KMeans GMM S-KM C-KM COPKM Agg DBSCAN DEC Proposed
dataset
1d_gauss 0.05 0.04 0.01 1.50 3.59 4.81 0.22 17.92 0.12
2d_gauss 0.02 0.14 0.01 5.89 4.80 1.40 0.07 12.17 0.11
6NewsgroupsUMAP10 0.02 0.11 0.01 2.25 1.52 0.42 0.09 6.76 0.12
MNIST_UMAP10 0.01 0.03 0.00 0.48 0.47 0.03 0.01 2.58 2.60
banknote 0.01 0.01 0.00 0.20 0.08 0.02 0.00 2.04 0.02
breast_cancer 0.01 0.10 0.00 0.05 0.12 0.00 0.01 0.85 0.01
cover_type 1.04 81.25 0.99 570.80 26.27
glass 0.00 0.01 0.00 0.07 0.64 0.00 0.00 0.37 0.79
ionosphere_UMAP10 0.00 0.00 0.00 0.02 0.02 0.00 0.00 0.59 0.01
iris 0.01 0.01 0.00 0.03 0.01 0.00 0.00 0.32 0.00
land_mines 0.01 0.03 0.00 0.11 0.17 0.00 0.00 0.58 0.70
pendigits 0.04 2.00 0.01 22.32 13.04 2.10 0.43 14.04 0.14
seeds 0.00 0.01 0.00 0.03 0.02 0.00 0.00 0.35 0.00
shuttle 0.17 0.52 0.03 8.16 0.55 46.73 0.77
wine 0.00 0.01 0.00 0.04 0.04 0.00 0.00 0.32 0.00
yeast 0.02 0.13 0.01 1.62 0.24 0.02 0.04 2.22 0.02
Table 7: Runtime (s) metric results
@@ 3.3.1 MNIST Handwritten Digits
This section presents the visual results of applying the proposed clustering algorithm to the MNIST digits dataset. Figure 15 depicts the application of UMAP to 2 dimensions using the cosine metric for visualisation. The plot with ground truth colouring reveals $10$ prominent clusters corresponding to each unique digit. However, several smaller groups are also observed indicating that the ground truth may not necessarily represent correct assignments since interpretation of the digits is sometimes ambiguous. The proposed clustering method is applied to the dataset after a UMAP transformation to 10 dimensions, and $5\%$ of the data ($90$ examples) is randomly sampled to serve as seeds for the clustering process. The clustering results are shown in Figure 16, where it achieves top tier performance according to the external validation metrics. However, a closer examination of the visual results also unveils an interesting finding: the algorithm not only successfully captures much of the primary clusters but also assigns points that may be outliers with respect to the clusters to the anomalous group. These anomalous parts of clusters or individual small clusters often represent either uniquely written digits, or slight variations within the same class, such as the digit $1$ with or without a horizontal baseline [^1] The interactive plots where points in the graph can be visualised are available in the GitHub project at https://github.com/M-Nassir/clustering.
Figure 15: The 2-dimensional UMAP of the MNIST digits dataset obtained from scikit-learn, using the cosine metric. Each data point is color-coded based on its corresponding class label obtained from the provided labels.
Figure 16: The results of applying the clustering algorithm to the MNIST digits data set obtained from sklearn after performing a UMAP transformation. The actual clustering process involved mapping the data to $10$ dimensions. However, for the purpose of visualisation, the results are presented in a 2-dimensional mapping.
@@
The 20 Newsgroups dataset in the scikit-learn library is often used as an example of clustering data in the text document space. It contains over $18,000$ news articles that are organised into $20$ categories. In this specific example, a subset of $6$ categories is used. These were selected for their distinctness, ease of visualisation and interpretation, because many of the labels have been reported to be noisy, and hence warrant possibly belonging to other classes or could equally belong to multiple classes. The UMAP reduction to $2$ dimensions is shown in Figure 17, where each example is coloured according to its ground truth category label. Notably, there is significant overlap between clusters, particularly in the central regions, and a small cluster consisting of examples from all categories can be observed on the far top-left of the plot.
Figure 17: A 2-dimensional UMAP plot of the Newsgroup data set is shown. Each point is coloured according to its ground truth label.
For clustering purposes, the data is mapped to $10$ dimensions and a $1\%$ random sample of labelled examples is used as seeds to guide the algorithm. The clustering results are shown in Figure 18 (displayed over the 2-dimensional mapping). The algorithm successfully recovers most of the six main clusters, achieving top-tier scores according to the external validation metrics. Moreover, the algorithm flags many examples located between clusters as anomalies, as well as the small dense cluster on the far top-left. This again highlights that the algorithm performs a clustering as opposed to a partitioning of the data points, where anomalies are also detected. Table 8 provides the top $10$ keywords for each of the six clusters obtained by the proposed method, along with those from the anomalous group. The top keywords align well with the category labels, while the anomalous group consists of seemingly random words that do not align well to any specific category.
Figure 18: The results of applying the proposed semi-supervised clustering algorithm on $6$ of the $20$ Newsgroup corpus dataset categories after a UMAP reduction to $10$ dimensions.
@@
Anomalies helmet, thanks, edu, just, like, list, mail, com, dog, wa
comp.windows.x use, display, application, program, motif, widget, thanks, file, server, window
rec.motorcycles rider, road, just, like, riding, dod, motorcycle, ride, wa, bike
rec.sport.hockey espn, season, playoff, play, year, player, hockey, wa, team, game
sci.crypt escrow, use, nsa, algorithm, phone, government, encryption, clipper, chip, key
soc.religion.christian christians, bible, people, sin, christ, christian, church, jesus, wa, god
talk.politics.guns just, weapon, government, law, did, right, fbi, people, wa, gun